A bicyclist rode into the county for 5 hours. In returning her speed was 5mi/h faster and her trip took 4 hours. What was her speed each way?
Answer by edjones(8007)

d=st d=dist., s=speed, t=time
st=st
Let s=speed going out. s+5=speed returning
5s=4(s+5)
5s=4s+20
5s-4s=4s-4s+20
s=20 mi/h, s+5=25 mi/h
Check:
20*5=25*4 true
Ed